# Memorization Over Reasoning? Exposing and Mitigating Verbatim Memorization in Large Language Models’ Character Understanding Evaluation

### v4

@@ 5 Hard Setting: Memorization-Controlled Prompting under Name Replacement
Furthermore, we aim to validate whether our prompting strategies can genuinely alter the model’s reasoning patterns under a hard setting, where memorization is explicitly restricted. Previous research has suggested that name replacement may impact model performance Yu et al. (2022). In this section, we confirm this hypothesis by demonstrating that name replacement significantly affects the model’s memorization ability and quantifying its impact. Then, we examine how prompting strategies influence LLMs’ reliance on memorization versus reasoning when memorization is constrained by name replacement.
@@ 5.1 Name Replacement Strategies and Their Impact on Memorization
  • Name Masking: Replacing all character names with placeholders (e.g., Sheldon → P0, Amy → P1).
  • Cross-Cultural Name Replacement: Substituting names with those from a different cultural background while maintaining gender consistency (e.g., SheldonJie Zhang, AmyXiaoling Wang). This tests whether models rely on specific name distributions for character inference.
  • Same-Cultural Name Replacement: Replacing names with other culturally consistent alternatives (e.g., SheldonAndrew, AmySally). This retains cultural familiarity while disrupting direct memorization.
@@
We design the Source Identification Task where the model is given a dialogue excerpt without source annotations and must predict the correct origin of the text. Let a dialogue segment be represented as $D=\{U_{1},U_{2},...,U_{n}\}$, where each utterance $U_{i}$ is part of a larger script from a specific source. The task requires the model to determine the correct source, $f_{\text{source}}(D)\rightarrow S$, where $S$ denotes the source work. We use the Fictional Works Memorization Benchmark Dataset described in Section 3.2.
@@
According to Figure 3, our results demonstrate that different name replacement strategies have a significant impact on the memorization ability of LLMs. Among them, Cross-Cultural Name Replacement consistently induces the strongest forgetting effect across all models, indicating that altering names to those from a different cultural background disrupts memorized associations more effectively than other replacement methods.
@@ 5.2 Memorization-Controlled Prompting under Name Replacement
Different name replacement strategies have shown a significant impact on the memorization ability of LLMs. Building on this, we apply the name replacement strategies introduced in Section 5.1 to evaluate the Character Guessing task using the different prompting strategies discussed in Section 4.2.
Under the Name Replacement condition, we provide brief character descriptions generated by GPT-4, aiming to enable the model to complete the Character Guessing task through character matching. The descriptions used are provided in the Appendix D. To assess the actual effectiveness of these descriptions, we conduct an ablation study.
@@
We present the results in Figure 4 and summarize our main findings as follows:
@@
    Strong Correlation Between Memorization and Performance. Model performance aligns closely with memorization levels, indicating a strong dependence on memorization for accurate predictions.
@@
    Minimal Impact of Character Descriptions . As an ablation study, we evaluate the effect of character descriptions under the Origin Script condition. Our results show that incorporating character descriptions does not significantly affect model performance, ruling out potential interference from inaccurate or misleading descriptions.
@@
    Verbatim Prompting Enhances Memorization. Across all three name replacement strategies, Verbatim prompting consistently improves model performance over the baseline, demonstrating that prompting can reinforce memorization and consequently boost accuracy.
@@
    Gist Prompting Reduces Reliance on Memorization. Performance under the Gist condition declines compared to the baseline, confirming that our Gist prompt effectively reduces the model’s reliance on memorization.
  5. 5.
    Name Replacement + Gist Prompting as an Effective Evaluation Strategy. We conclude that combining Cross-Culture Name Replacement with Gist prompting provides a reasonable and effective method for evaluating a model’s true capabilities while minimizing reliance on memorization.
@@ 6 Evaluating Character Understanding Tasks Under Memorization Constraints
In this section, we evaluate different types of character understanding tasks using the Gist Prompting with Cross Culture Name Replacement approach. We first introduce the design principles for Gist Prompts across various task types, followed by an analysis of the results obtained from three different LLMs.
@@ 6.1 Benchmark Tasks Description
We define the five character understanding tasks being benchmarked and provide an illustrative example.
Let a dialogue segment be represented as $D=\{U_{1},U_{2},...,U_{n}\}$, where $U_{i}$ is an utterance in the conversation. Each utterance $U_{i}$ is associated with a speaker label $S_{i}$, and additional annotations depending on the task.
  • Coreference Resolution: Cluster character mentions $C=\{c_{1},c_{2},...,c_{n}\}$ into groups referring to the same entity, $f_{\text{coref}}(C)\rightarrow\{\text{clusters of }c_{i}\}$.
  • Personality Understanding: Classify a character’s personality trait $P$ given dialogue context $D$, $f_{\text{personality}}(D,S_{i})\rightarrow P$.
  • Role Detection: Determine whether a character plays a key role (e.g., perpetrator in a crime narrative), $f_{\text{role}}(S_{i})\rightarrow\{0,1\}$.
  • Open-Domain Question Answering: Extract the span $A$ in $D$ that answers a given question $Q$, $f_{\text{QA}}(D,Q)\rightarrow A$.
  • Summarization: Generate a summary $S$ that captures key events and relationships, $f_{\text{summ}}(D)\rightarrow S$.
@@ 6.2 Gist Prompts Across Character Understanding Tasks
Detailed gist prompts are included in Appendix H.
@@ 6.3 Results and Key Findings
Our main results and findings are summarized as follows:
@@
    Significant Performance Drop with Name Replacement: Across all tasks, we observe a substantial decline in performance when applying the name replacement (NR) method in Table 6.2. This result indicates that reducing explicit memorization has a strong impact on model performance, confirming our hypothesis that existing character understanding tasks rely heavily on memorization. Moreover, it highlights the potential inflation of benchmark scores under standard evaluation settings.
@@
    Further Degradation with Gist Prompting: The additional performance drop induced by Gist prompting demonstrates its effectiveness across various character understanding tasks. This suggests that Gist prompting can further encourage models to rely less on memorization, reinforcing our findings on the role of memorization in these tasks.
@@
    Memorization Can Be Detrimental to Character Understanding: Interestingly, we observe performance gains in certain cases under the combined NR+Gist setting. Our error case analysis reveals that these improvements stem from resolving knowledge conflicts—instances where the original memorization-based responses misled the model, whereas reducing memorization mitigated such errors, leading to better predictions.
  4. 4.
    Consistent Trends Across Models: As shown in Figure 5, despite variations in absolute performance, different models exhibit similar trends in performance degradation across all tasks. This consistency suggests that various task designs interact similarly with memorization, reinforcing the broader implications of our findings.
Figure 5: Performance degradation across LLMs due to the effects of GIST+NR perturbations.
@@ 7 Conclusion
In this work, we introduce the concept of gist memory and design prompts for six types of character understanding tasks, enabling models to reduce reliance on memorization and enhance reasoning-based character interpretation. Our approach provides a systematic framework to evaluate the extent to which different character understanding tasks depend on memorization, offering a more accurate assessment of model performance by disentangling genuine reasoning from memorization-driven responses. These findings highlight the limitations of current benchmarks and provide insights for designing more robust character understanding evaluations.
@@ 8 Limitations
Our study can be extended by future research in a number of ways. First, though we ground and motivate our work through “verbatim” and “gist” memory, we note that any apparent “reasoning” of AI systems may differ fundamentally from those of humans, and we did not pursue human studies to examine this in detail. Further investigation is needed to explore this topic.
@@ 9 Ethics
Our study probes the degree to which GPT-4o have memorized fictional works, and its impact on character understanding tasks. This work uses the OpenAI API on experiments above, and at no point do we access, or attempt to access, the true training data behind these models, or any underlying components of the systems. Risks The several datasets and fictional works scripts in our experiment are sourced from publicly available sources. However, we cannot guarantee that they are devoid of socially harmful or toxic language. We use ChatGPT  [^7] https://chatgpt.com/ to correct grammatical errors in this paper.
@@ Appendix A Example of Script-based Character Understanding Task
In Table 4 we show a full example of the character understanding task. In this example, we have automatically replaced the individual character names with placeholders like P0 and P1.
@@
Appendix B Example for Verbatim Memory
B.1 Source Prediction
@@
Appendix C Extended Dataset Analysis
@@
Table 6: Overview of 14 tasks in the character understanding benchmark.
Appendix D Prompts for GPT4
for personality traits extraction ’you are a psychology research assistant designed to help analysis character personalities according to conversations.’ conversation ’Here ends the conversation.Give me a character description for each main character according to this conversation.’ Character Description Here is a character description:Monica: Responsible, caring, and organized. Assertive and confident in her actions.Chandler: Witty and self-deprecating with an approachable sense of humor. Exhibits insecurity and anxiety in his dialogues, making references to uncomfortable situations and questioning his own actions.Rachel: Spontaneous and open to change, she takes risks and is adaptable. She is also reliant on her relationships with others, signifying her dependency and need for support.’ ’Joey: Energetic, extroverted, and casual. Lacks the sensitivity of others feelings at times but genuinely care about friends.Phoebe: Quirky, eccentric, and a free spirit. Her train of thought tends to lean toward the unusual and bizarre. However, she is also compassionate and caring.Ross: Insecure and somewhat neurotic and vulnerable. His behavior is indicative of someone who is going through emotional turmoil.’
Appendix E Name Replacement and Character Comparison
E.1 Cross-Language Uncommon Name Replacement
@@
E.2 Same-Language Replacement
@@
E.3 Anonymized Speaker Replacement
@@
E.4 Control Test
Example Here is a dialogue between 2 people: Sheldon : (now lifts her head) phoebe!! Amy : you know it’s me? Sheldon : for like a half an hour! man, you can lie about sweden! Amy : how can you come here? Sheldon : how could you not tell me you worked here? Amy : i don’t have to tell you everything!P2: We got an eight-point-two. Trust me, you’re happy.
Prompt The following is a dialogue between the same two people. Please identify who they are:
P0 : (singing) ïpan stripan, glupi glabi! änd that is the swedish national anthem! thank you for asking! (looks annoyed) P1 : wow, ikea… what a rich culture. uhm, you know what? i have a friend who is a masseuse. P0 : oh, ja! ja! P1 : yah! she’s… uhm… not very good though… (phoebe looks devastated)
Table 8: An example for the control test, where in the example,the script is from Friends but speakers are annoted as Amy and Sheldon(from The Big Bang Theory).
Appendix F Verbatim avoid Reasoning
Figure 6: In a personality understanding task where the scenario involves ’Joey returned the lost wallet to its rightful owner, Monica,’ the question posed might be, ’According to this action, his personality should be: ?’ The human reasoning process would draw on contextual evidence, concluding that returning lost property demonstrates honesty. In contrast, the LLM might conclude, ’Joey and Monica are from the TV show Friends, and Joey is characterized as an honest character in this show,’ thus relying heavily on memorized information rather than genuine reasoning.
Appendix G Benchmark Datasets
@@
  • TVShow Character Guessing: Based on the TVSHOWGUESS dataset, which contains scripts from five popular TV series: Friends, The Big Bang Theory, The Office, Frasier, and Gilmore Girls Sang et al. (2022c), this is a multi-choice classification task. The dialogue is anonymized using placeholder identifiers (e.g., P1, P2), and the model must predict the correct speaker for each line, testing its ability to identify characters based on context.
@@
Appendix H Gist Prompting Strategies Across Tasks
@@
H.1 Task-Specific Gist Prompts

### v5

@@ 5 Hard Setting: Memorization-Controlled Prompting under Name Replacement
To effectively suppress verbatim memorization and induce gist-style reasoning, we design a stronger intervention: replacing character names to explicitly remove memorized surface cues. While prior work uses name replacement to improve model robustness Yu et al. (2022), we take this idea further by systematically testing whether such name-based interventions are strong enough to disrupt verbatim memorization and induce gist-based reasoning.
@@ 5.1 Name Replacement Strategies and Their Impact on Memorization
  • Name Masking: Replacing character names with placeholders (e.g., Sheldon, Amy → P0, P1).
  • Cross-Cultural Name Replacement: Substituting names with those from a different cultural background (e.g.English → Chinese) while maintaining gender consistency (e.g., Sheldon, AmyJie Zhang, Xiaoling WangSun et al. (2024). This tests whether models rely on specific name distributions for character inference.
  • Same-Cultural Name Replacement: Replacing names with other culturally consistent alternatives (e.g., Sheldon, AmyAndrew, Sally). This retains cultural familiarity while disrupting direct memorization by name replacements.
@@
To directly quantify how much verbatim memorization models rely on, we design a Source Identification Task. This task closely resembles established memorization evaluation methods, such as source attribution or textual cloze tests Chang et al. (2023b). Specifically, given a dialogue excerpt $D=\{U_{1},U_{2},...,U_{n}\}$ without source annotations, the model must identify the correct origin of the text, predicting $f_{\text{source}}(D)\rightarrow S$, where $S$ is the specific fictional source.
The rationale behind this design is straightforward: if a model has memorized verbatim text from pretraining, it should accurately identify the source based purely on surface-level textual recall. Conversely, if name replacement or prompting disrupts memorization effectively, we expect a significant decrease in source-identification accuracy. Thus, observing substantial performance degradation would strongly support the claim that models’ prior accuracy was driven largely by verbatim memorization rather than reasoning. We use the CharScript: Fictional Works Memorization Benchmark Dataset described in Section 3.2 to rigorously evaluate this effect.
@@
According to Figure 3, our results demonstrate that different name replacement strategies have a significant impact on the memorization ability of LLMs. Among them, Cross-Cultural Name Replacement consistently induces the strongest forgetting effect across all models, indicating that altering names to those from a different cultural background disrupts memorized associations more effectively than other replacement methods. This suggests that cultural distance may play a crucial role in weakening memorized identity-linking patterns in models.
@@ 5.2 Memorization-Controlled Prompting under Name Replacement
Different name replacement strategies have shown a significant impact on the memorization ability of LLMs. Building on this, we apply the name replacement strategies introduced in Section 5.1 to evaluate the Character Guessing tasks, using the verbatim and gist prompting strategies we have discussed in Section 4.2.
Under the Name Replacement condition, we provide brief character descriptions generated by GPT-4, aiming to enable the model to complete the Character Guessing task through character matching. The descriptions used are provided in the Appendix C. To assess the actual effectiveness of these descriptions, we conduct an ablation study.
@@
We present the core results in Figure 4 and summarize four key findings:
@@
    Performance Declines with Reduced Memorization. Accuracy drops as surface cues are removed, revealing the model’s strong reliance on verbatim memorization.
@@
    Character Descriptions Have Minimal Impact. Adding character descriptions does not notably change performance, suggesting limited influence on model decisions.
@@
    Prompting Style Matters. Verbatim prompts boost performance by reinforcing memorization, while Gist prompts suppress it and lead to lower accuracy.
@@
    Name Replacement + Gist Prompting Is Effective. This combination best reduces memorization reliance and reveals the model’s true reasoning ability.
@@ 6 Evaluating Character Understanding Tasks Under Memorization Constraints
Sections 4 and 5 provided clear empirical evidence supporting our hypothesis: LLMs’ strong performance on character understanding tasks significantly relies on verbatim memorization. To further examine the generalizability of our approach and assess how memorization affects performance across diverse task types, we now extend the combination of Gist Prompting and Cross-Cultural Name Replacement to five additional character understanding benchmarks.
@@ 6.1 Benchmark Tasks Description
We benchmark five types of character understanding tasks spanning coreference, personality inference, role detection, QA, and summarization. Each task evaluates the model’s ability to extract or infer character-related information from dialogue.
@@ 6.2 Gist Prompts Across Character Understanding Tasks
Detailed prompts in Appendix F.
@@ 6.3 Results and Key Findings
Our evaluation reinforces our conclusions regarding memorization versus reasoning, with three key findings:
@@
    Significant Performance Drop with Reduced Memorization: As shown in Table 6.2, applying the name replacement (NR) method leads to a substantial performance decline across all tasks. This confirms our hypothesis that current character understanding benchmarks are heavily reliant on explicit memorization, potentially inflating standard evaluation scores. Furthermore, prompting models to use gist reasoning induces an additional performance drop, underscoring the effectiveness of gist prompting in mitigating reliance on surface-level memorization.
@@
    Performance Gaps across Task Types (Motivation vs. Action Prediction): We observe clear differences in memorization effects between action-oriented tasks (e.g., Character Guessing, CSI Role Detection) and motivation-oriented tasks (e.g., Personality Understanding). Action prediction tasks typically require models to output discrete labels—such as a character name or role—making it easy for models to memorize associations between surface cues and labels. For instance, phrases consistently linked to characters like Sheldon (“I’m not crazy, my mother had me tested”) are easily memorized, but become brittle under name replacement. In contrast, motivation prediction tasks involve deeper contextual inference about characters’ internal states and intentions, rendering surface-level memorization less effective, thus promoting genuine reasoning.
@@
    Memorization Can Be Detrimental: Interestingly, certain tasks show performance improvements under the NR+Gist setting. Our error analysis reveals that reducing memorization resolves knowledge conflicts—cases where memorized but incorrect associations previously misled models, indicating that excessive memorization can actively impair true character understanding.
These findings are consistent across multiple models (Figure 5), suggesting generalizability and broader implications of memorization biases in current benchmarks.
Figure 5: Performance degradation across LLMs due to GIST+NR. Action prediction tasks (e.g., Guessing, CSI) show greater drops than motivation prediction tasks (e.g., PERSONET, FriendsQA), reflecting their differential reliance on memorization.
@@ 7 Conclusion
In this work, we introduce the distinction between verbatim memorization and gist memorization, and systematically demonstrate that many existing character understanding evaluations rely heavily on verbatim memorization, potentially inflating perceived reasoning capabilities of LLMs. To rigorously test and mitigate this reliance, we propose a novel, graded intervention framework combining Gist Prompting and Cross-Cultural Name Replacement. Our experimental results across six diverse character-centric benchmarks clearly show that reducing verbatim memorization cues significantly affects model performance, highlighting the importance of explicitly disentangling memorization from genuine reasoning.
Our findings emphasize the critical need for developing memorization-aware evaluations, particularly in domains like character understanding where memorization effects are pronounced. Broadly, our methodology offers generalizable insights applicable to a wide range of inference tasks where accurately distinguishing between memorization and reasoning is essential.
@@ 8 Limitations
Our study can be extended by future research in a number of ways. First, though we ground and motivate our work through “verbatim” and “gist” memorization, we note that any apparent “reasoning” of AI systems may differ fundamentally from those of humans, and we did not pursue human studies to examine this in detail. Further investigation is needed to explore this topic.
@@
Thirdly, While our interventions reveal strong trends, we acknowledge that results may be sensitive to the surface form of prompts. Future work could complement our findings with uncertainty-based analyses such as log probability comparisons.
@@ 9 Ethics
Our study probes the degree to which GPT-4o have memorized fictional works, and its impact on character understanding tasks. This work uses the OpenAI API on experiments above, and at no point do we access, or attempt to access, the true training data behind these models, or any underlying components of the systems. Risks The several datasets and fictional work scripts in our experiment are sourced from publicly available sources. However, we cannot guarantee that they are free from socially harmful or toxic language. We use ChatGPT  [^1] https://chatgpt.com/ to correct grammatical errors in this paper.
@@ Appendix A Example of Script-based Character Understanding Task
In Table˜4 we show a full example of the character understanding task. In this example, we have automatically replaced the individual character names with placeholders like P0 and P1.
@@
Appendix B Extended Dataset Analysis
@@
Table 6: Overview of 8 more tasks in the character understanding benchmark.
Appendix C Prompts for GPT4
The following is the prompt we sent to GPT4: For personality traits extraction ’you are a psychology research assistant designed to help analysis character personalities according to conversations.’ conversation ’Here ends the conversation.Give me a character description for each main character according to this conversation.’
Character Description Here is a character description:Monica: Responsible, caring, and organized. Assertive and confident in her actions.Chandler: Witty and self-deprecating with an approachable sense of humor. Exhibits insecurity and anxiety in his dialogues, making references to uncomfortable situations and questioning his own actions.Rachel: Spontaneous and open to change, she takes risks and is adaptable. She is also reliant on her relationships with others, signifying her dependency and need for support.’ ’Joey: Energetic, extroverted, and casual. Lacks the sensitivity of others feelings at times but genuinely care about friends.Phoebe: Quirky, eccentric, and a free spirit. Her train of thought tends to lean toward the unusual and bizarre. However, she is also compassionate and caring.Ross: Insecure and somewhat neurotic and vulnerable. His behavior is indicative of someone who is going through emotional turmoil.’
Appendix D Name Replacement and Character Comparison
We also provide partial name replacement dictionary in table  7.
D.1 Cross-Language Uncommon Name Replacement
@@
D.2 Same-Language Replacement
@@
D.3 Anonymized Speaker Replacement
@@
Appendix E Benchmark Datasets
@@
  • TVShow Character Guessing: Based on the TVSHOWGUESS dataset, which contains scripts from five popular TV series: Friends, The Big Bang Theory, The Office, Frasier, and Gilmore Girls Sang et al. (2022b), this is a multi-choice classification task. The dialogue is anonymized using placeholder identifiers (e.g., P1, P2), and the model must predict the correct speaker for each line, testing its ability to identify characters based on context.
@@
Appendix F Gist Prompting Strategies Across Tasks
@@
F.1 Task-Specific Gist Prompts